-4x - 5y - 8z = 268
-10x + 7y + 6z = -786
-12x + 9y - 2z = -548


I need to solve this using any method, but id like to see the steps and the correct answer. I'm not sure where to start with this one, im just lost. Thank you so much for all your help.
I tried to do this on my own, i thought i should multiply the first equation by 3 and add to the 3rd equation,but im not sure what to do next, or if that was even where to start.

Answer by scott8148(5880)

looking at the values of the coefficients (numbers), the numbers for the z terms seem the easiest for eliminating

multiply 3rd eqn by 3, add to 2nd eqn __ -46x+34y=-2430 __ -23x+17y=1215
__ multiplying by 41 __ -943x+697y=-49815 __ (A)

multiply 3rd eqn by 4, subtract from 1st eqn __ 44x-41y=2460
__ multiplying by 17 __ 748x-697y=41820 __ (B)

adding eqns (A) and (B) __ -195x=-7995 __ x=41

substituting __ 44(41)-41y=2460 __ y=-16

substituting __ -4(41)-5(-16)-8z=268 __ z=-44